OpenStudy (anonymous):

The two figures are similar. The area of the smaller trapezoid is 181. What is the area of the bigger trapezoid? The ratio of the smaller trapezoid to the bigger trapezoid is 14:26. a) 7 b) 676 c) 624 d)196

OpenStudy (misssunshinexxoxo):

Any graph?

OpenStudy (kropot72):

The area of a trapezoid is given by \[\large Area=\frac{1}{2}h(a+b)\] where the altitude is h and the parallel sides are a and b. Let the scale factor 14 : 26 be represented by s. Then the area of the larger trapezoid will be \[\large Area _{\lg}=\frac{1}{2}hs(as+bs)=\frac{1}{2}hs^{2}(a+b)\] Now can you see that the area of the larger trapezoid is found by multiplying the are of the smaller by \[\large (\frac{26}{14})^{2}\]

OpenStudy (kropot72):

@kristinalgarcia Are you there?

OpenStudy (anonymous):

so it's 676?

OpenStudy (kropot72):

The area of the larger trapezoid is given by \[\large 181\times (\frac{26}{14})^{2}=you\ can\ calculate\]

OpenStudy (kropot72):

You need to redo the above calculation.

OpenStudy (anonymous):

oh so it's 624 then!

OpenStudy (kropot72):

Yes, you are correct.

OpenStudy (anonymous):

Thank you so much!!!